Find the differential equations of the following equations by integrating factors by inspection. Show complete solution.


xdy - ydx = x^4 ydy + x^3 y^2dx 


1
Expert's answer
2021-10-12T09:01:40-0400

"\displaystyle
xdy -ydx = x^4ydy + x^3y^2dx\\
=xdy -ydx - x^4ydy-x^3y^2dx=0\\
=xdy - ydx - x^3y(xdy+ydx)=0\\
\text{Divide through by $x^2$ to obtain}\\
\frac{xdy - ydx}{x^2}-xyd(xy)=0\\
=d(\frac{y}{x})-xyd(xy)=0 \qquad (1)\\
\text{Integrating (1) we have}\\
\frac{y}{x}-\frac{1}{2}(xy)^2=c"